Please help. Not sure what book this comes from.
Solve. W=Fg/a(h1-h0) for h0. Not sure how to indicate that the 1 and 0 are lower than the h. I'm not sure what to do with the Fg/a, but I came up with h1=h0/h0-1.
THank you!

Answer by stanbon(48558)

Solve. W=Fg/a(h1-ho) for ho
It is hard to tell if your (h1-ho) is in the numerator or
the denominator. I'll assume it is the numerator as follows:
W=(Fg/a)(h1-ho)
Multiply both sides by a to get:
Wa=F(h1-ho)
Divide both sides by F to get:
(Wa/F)=h1-ho
Subtract h1 from both sides to get:
(Wa/F)-h1=-ho
Multiply both sides by -1 to get:
ho=-(Wa/F)+h1
Cheers,
Stan H.